In testing a engine, various mixtures of gasoline and methanol are being tried. How much of a 90% gasoline mixture and a 75% gasoline mixture are needed for 1200L of an 85% gasoline mixture?
Answer by stanbon(75887)

How much of a 90% gasoline mixture and a 75% gasoline mixture are needed for 1200L of an 85% gasoline mixture?
------
Equation:
gas + gas = gas
0.90x + 0.75(1200-x) = 0.85*1200
---
Multiply thru by 100 to get:
90x + 75*1200 - 75x = 85*1200
---
15x = 10*1200
x = (2/3)1200
x = 800 L (amt. of 90% mixture needed)
---
1200-x = 400 L (amt of 75% mixture needed)
============================================
Cheers,
Stan H.
===========